In 5 years Tom's age will be twice what it was 6 years ago. Write an equation for how old Tom is now?

Answer by Naruto4085(1)

Question: In 5 years Tom's age will be twice what it was 6 years ago. Write an equation for how old Tom is now?
Present Date: 2015
In Five Years- 2020
So that means in 2020, Tom's age will be twice what it was 6 yrs ago. Six years ago is in 2009.
Let x= Tom's age Equation- 5+x=2(x-6)
Lets solve for x
5+x=2(x-6) Rewrite Your Equation
5+x=2x-12 Distribute
5=X-12 Subtract x on both sides
x= 17 Add 12 on both sides
Tom is 17 years old.

The Beast